• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jack

    Log entry time 19:20:35 on June 29, 2010

    Entry number 329160

    keyword=Right VDCs gas flow

    The gas flow to the VDCs on the Right Arm was restarted at about 6PM. New bubblers have been installed on the inputs and outputs. The rotameters had the black beads at about 40, a little higher than they were set previously. Some air was certainly introduced into the system. So, allow plenty of time for the chambers to flush out before high voltage is applied. The high voltage cables are disconnected at the aluminum boxes on the VDC frame, the box that has the three cables going to the VDC.



    A copy of this log entry has been emailed to: bogdanw@jlab.org